A 10-kg projectile is fired vertically upward from the ground, with an initial velocity of 50 m/s, Fig. 13-7a. Determine the maximum height to which it will travel if (a) atmospheric resistance is neglected; and (b) atmospheric resistance is measured as Fo=(0.01v) N, where v is the speed of the projectile at any instant, measured in m/s.

1
Expert's answer
2022-03-28T11:02:56-0400

"h_1=\frac{v^2}{2g}=127.4~m,"

"h_2=\frac{v^2}{2(g+\frac Fm)}=126.8~m."